OpenStudy (anonymous):

Help me again please . . I honestly don't get this 1/2 = 32^x

OpenStudy (marigirl):

When the power of an equation is unknown we need to apply Index laws (which includes logarithmic laws) \[\log (0.5)=\log (32)^x\] \[\log (0.5)=x \log (32)\] then solve for x \[x=\log (0.5)/\log(32)\]

OpenStudy (anonymous):

just remember \[b ^{y}=x\] and \[\log _{b}x=y\]

OpenStudy (anonymous):

Thanks guys !!